cos(x+pi/2)=-sin x

Answer by math-vortex(648)

Hi there,

PROBLEM:
Prove that the identity is true.
cos(x+pi/2)=-sin (x)

Solution:
Transform the left side of the identity until is is the same as the right side.
cos(x + pi/2) = -sin (x)
cos(pi/2 - (-x)) = -sin (x)

Apply the co-function property: [cos(pi/2 + x) = sin (x)]
sin (-x) = -sin (x)

Apply the even/odd formula: [sin (-x) = - sin (x)]
-sin (x) = - sin (x)

That's it.

Mrs. Figgy
math.in.the.vortex@gmail.com